Show that the circular cylinder S = {(x, y, z) ∈R^3 : y^2+z^2 = 1}can be covered by a single regular surface patch, and hence is a surface.
1
Expert's answer
2020-10-19T16:01:14-0400
"Solution"

A surface patch can be given by "(x,y)\mapsto (\frac{x}{r},\frac{y}{r}, tan (r-\pi/2))."

However we must be careful on how we define the open set that this is defined over

Here "\{x,y|\sqrt {x^2+y^2} \in (0,\pi)\}."

Also, ."r=+\sqrt{x^2+y^2}".

Clearly, the map is smooth since "r\neq 0." Also "r<\pi."

Hence "tan(r-\pi/2)" tan is well defined.

For surjection



Let "(x,y,z)" be a given point on the cylinder. Since tan is continuous and injective on the range "(-\pi/2,\pi/2)" and unbounded its surjective, and hence we get "r" such that "tan(r-\pi/2)=z." So take "a=yr, b=zr." Now "\sqrt{y^2+z^2}=1\Rightarrow y,z\leq 1""\sqrt{y^2+z^2}=1\Rightarrow y,z\leq 1" . Hence "a,b\leq r<\pi". Hence the pre-image lies in our given domain.